Find the integral of: x / sqrt(4x^2 + 6x) dx with limits of range increasing from -1/2e to 1/2e
1
Expert's answer
2021-01-26T03:19:03-0500

Set "I = \int \frac {x}{\sqrt{(4x^{2} +6x)}} dx"

Completing the square in the denominator, we have that

"I = \int \frac {2x}{\sqrt{(4x + 3)^{2} -9}} dx"

Applying the substitution "4x +3 = 3 cosh \theta"

We have that

"I = \int \frac{2}{4} (\frac{3cosh\theta - 3}{\sqrt{9cosh^{2}\theta - 9}})(\frac{3}{4}sinh\theta d\theta)"

By simplifying, we have

"I = \frac{3}{8} \int (cosh\theta -1) d\theta"

Integrating term by term

"I = \frac{3}{8}(sinh\theta - \theta) + C"

Reverse the substitution

"I = \frac{3}{8}[(\frac{1}{3})\sqrt{(4x+3)^2-9} - \cosh^{-1}(\frac{4x + 3}{3})] + C"

Applying the upper and lower limits to the value of the definite integral

"\int_ {\frac{-1}{2e}}^{\frac{1}{2e}} \frac {x}{\sqrt{(4x^{2} +6x)}} dx" is approximately "0.3504 + 0.7257i"

The reason why we have a complex solution is because of the lower limit of the integration.